The period of a particle executing SHM is 8 s. At t= 0 it is at the mean position. The ratio of the distances covered by the particle in the 1st second to the 2nd second is

Grade:12th pass

2 Answers

Arun
5 years ago
let the equation of motion is X =Asin(2px/8) in first second X1 =Asin(p/4) in 2 sec X2 =Asin(p/2) so 2nd sec distance covered = X2 -X1 x1 /(X2 -X1) = Asin(p/4)/[Asin(p/2) -Asin(p/4)] =1/(v2 -1
AKSHANSH PAREEK
5 years ago

Y¹= A sin w = A sin 2π\8 = A /√2 

Distance when t=2 sec is

Y²= A sin w= A sin π\2= A 

Distance covered in 2sec is A-A/√2 

Y¹/Y²= (A/√2) / (A-A/√2) 

= 1/√2-1

please approve